# Spatial Reasoning with Vision-Language Models in Ego-Centric Multi-View Scenes

### v1

@@ A.3 Benchmarking Generalist VLMs: Further Results
A.4 Inference Time Analysis
Table 11 shows the inference time analysis of Ego3D-VLM compared to baseline models. We report End-to-End Latency (E2E Lat.) in seconds and Peak Memory in GB. For the E2E Lat., we measure the average end-to-end inference of models on 50 samples of Ego3D-Bench and for the peak memory we report the peak memory usage during inference on the same samples. Experiments are performed on NVIDIA H20 GPUs using flash-attention-2. The memory and latency overhead of Ego3D-VLM over InternVL3-78B is 0.6% and 31% , respectively. The main reason for the latency overhead is that the model reasons more when cognitive map is provided in the Ego3D-VLM.
@@ A.4 Inference Time Analysis
A.5 Robustness of External Tools
To isolate tool performance, we included an ablation with ground-truth cognitive maps (Table 6), which improve the results from 60.1% to 79.4% accuracy. This confirms that, there is still a gap between the accuracy of such imperfect tools and ground-truth REC/depth. However, compared to the baseline, our solution still achieves more robust results even in challenging conditions such as low brightness, motion blur, and occlusion. To support this claim, we simulated these conditions and re-ran more specific experiments shown in the Table 12.
@@ A.5 Robustness of External Tools
Although G-DINO and Depth-Anything used in Ego3D-VLM are among the best tools for REC and depth estimation, our solution is orthogonal to any such tools, and has no limitation in this regard.
A.6 Qualitative Results
Figure 9-19 demonstrate example responses of InternVL3-78B and Ego3D-VLMInternVL3-78B on all categories of Ego3D-Bench. As seen, Ego3D-Bench enhances the spatial reasoning ability of the baseline by providing the textual cognitive map.
A.7 Benchmarking the Thinking of Ego3D-Bench as an Open-QA
Thinking of Obj. Abs. Dist. Thinking of Ego Abs. Dist.
$\uparrow GPT-4o Score (0-10)\uparrow$
InternVL3-8B 1.9 2.0
Ego3D-VLM InternVL3-8B 4.7 5.2
InternVL3-14B 2.3 2.4
Ego3D-VLM InternVL3-14B 5.7 5.0
InternVL3-38B 2.1 2.5
Ego3D-VLM InternVL3-38B 7.1 3.7
InternVL3-78B 2.1 1.7
Ego3D-VLM InternVL3-78B 7.0 6.8
A.8 Qualitative Results
Figure 9-19 demonstrate example responses of InternVL3-78B and Ego3D-VLMInternVL3-78B on all categories of Ego3D-Bench. As seen, Ego3D-Bench enhances the spatial reasoning ability of the baseline by providing the textual cognitive map.
@@ A.9 Limitations and Future Work
Another limitation lies in the accuracy of metric depth estimation in outdoor environments. To mitigate this, we proposed a relational scaling approach based on common-sense object sizes. However, this method is inherently approximate and not fully reliable. Future work should aim to address these issues by enhancing the depth understanding of VLMs or by improving metric depth estimation in outdoor settings. Better integration of spatial reasoning and scene geometry into VLMs could further improve performance in complex 3D environments.

### v2

@@ A.4 Benchmarking Generalist VLMs: Further Results
A.5 Inference Time Analysis
Table 11 shows the inference time analysis of Ego3D-VLM compared to baseline models. We report End-to-End Latency (E2E Lat.) in seconds and Peak Memory in GB. For the E2E Lat., we measure the average end-to-end inference of models on 50 samples of Ego3D-Bench and for the peak memory we report the peak memory usage during inference on the same samples. Experiments are performed using flash-attention-2. The memory and latency overhead of Ego3D-VLM over InternVL3-78B is 0.6% and 31% , respectively. The main reason for the latency overhead is that the model reasons more when cognitive map is provided in the Ego3D-VLM. In order to make the inference more efficient and deal with the latency overhead of the VLMs, post-training techniques such as quantization [15], token pruning [1], or knowledge distillation [13] can be used.
@@ A.5 Inference Time Analysis
A.6 Robustness of Ego3D-VLM in Challenging Conditions
To isolate tool performance, we included an ablation with ground-truth cognitive maps in the main body of the paper (Table 6), which improve the results from 60.1% to 79.4% accuracy. This confirms that, there is still a gap between the accuracy of such imperfect tools and ground-truth REC/depth. However, compared to the baseline, our solution still achieves more robust results even in challenging conditions such as low brightness, motion blur, and occlusion. To support this claim, we simulated these conditions and re-ran more specific experiments shown in the Table 12. Although G-DINO and Depth-Anything used in Ego3D-VLM are among the best tools for REC and depth estimation, our solution is orthogonal to any such tools, and has no limitation in this regard.
@@ A.6 Robustness of Ego3D-VLM in Challenging Conditions
A.7 Results of different Source Datasets
NuScenes Waymo Argoverse
[…]
Closed-source Models
GPT-4o 60.9 21.2 60.0 19.0 59.4 17.1
Ego3D-VLM GPT-4o 73.5 8.5 73.1 7.5 72.7 6.2
Gemini-1.5-Pro 59.3 26.4 55.6 12.7 57.5 19.5
Ego3D-VLM Gemini-1.5-Pro 77.1 8.1 68.9 6.3 73.0 7.1
Qwen2.5 Family
Qwen2.5-3B 39.7 33.4 38.6 31.5 38.9 31.2
Ego3D-VLM Qwen2.5-3B 44.7 13.0 40.3 12.1 48.1 12.6
Qwen2.5-7B 40.9 36.0 40.6 24.7 41.6 30.1
Ego3D-VLM Qwen2.5-7B 56.4 11.3 51.9 8.3 54.4 8.7
Qwen2.5-32B 56.6 17.5 56.0 11.0 59.2 18.7
Ego3D-VLM Qwen2.5-32B 64.3 17.1 64.8 10.6 64.0 12.1
Qwen2.5-72B 58.2 21.2 57.7 11.1 58.2 16.2
Ego3D-VLM Qwen2.5-72B 74.4 8.5 64.4 6.4 69.5 7.6
InternVL3 Family
InternVL3-8B 42.3 28.3 44.0 21.7 42.9 31.5
Ego3D-VLM InternVL3-8B 62.0 10.5 57.1 6.5 61.3 6.7
InternVL3-14B 51.7 20.4 50.2 15.8 53.3 16.6
Ego3D-VLM InternVL3-14B 69.0 8.8 63.7 7.5 67.0 6.3
InternVL3-38B 51.2 31.4 50.5 19.4 53.4 25.4
Ego3D-VLM InternVL3-38B 71.7 8.7 64.3 8.2 67.0 6.3
InternVL3-78B 60.0 16.7 59.7 10.7 60.1 13.7
Ego3D-VLM InternVL3-78B 76.5 8.3 67.0 6.5 71.0 7.5
Ovis2 Family
Ovis2-4B 41.2 29.2 43.0 18.1 40.9 22.5
Ego3D-VLM Ovis2-4B 47.3 9.8 47.5 7.1 49.7 8.3
Ovis2-8B 41.5 21.4 41.9 20.7 41.0 20.2
Ego3D-VLM Ovis2-8B 55.6 9.4 55.1 6.8 57.9 7.0
Ovis2-16B 48.6 16.2 45.5 11.9 46.8 12.9
Ego3D-VLM Ovis2-16B 62.6 8.5 57.9 7.15 62.2 6.5
A.8 Qualitative Results and Prompts Structure
Figures 9-19 demonstrate example responses of InternVL3-78B and Ego3D-VLMInternVL3-78B on all categories of Ego3D-Bench. As seen, Ego3D-Bench enhances the spatial reasoning ability of the baseline by providing the textual cognitive map.
@@ A.9 Limitations and Future Work
Another limitation lies in the accuracy of metric depth estimation in outdoor environments. To mitigate this, we proposed a relational scaling approach based on common-sense object sizes. However, this method is inherently approximate and not fully reliable. Future work should aim to address these issues by enhancing the depth understanding of VLMs or by improving metric depth estimation in outdoor settings. Better integration of spatial reasoning and scene geometry into VLMs could further improve performance in complex 3D environments. Another major limitation in spatial reasoning datasets and benchmarks is the availability of ground-truth 3D. Previous work has shown the effectiveness of neural radiance fields in generating photo-realistic 3D data [14]. Future work should investigate this direction for extending spatial reasoning dataset without having ground-truth 3D annotations.
A.10 Statement on LLMs Assistance
We declare that some portions of this document have been lightly refined using Large Language Models (e.g., ChatGPT) to enhance clarity and polish. All substantive content and ideas remain entirely our own.